# Tiny is not small enough: High quality, low-resource facial animation models through hybrid knowledge distillation

### v1

@@ Abstract.
submissionid: 1216journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3730929ccs: Computing methodologies Neural networksccs: Computing methodologies Procedural animation
@@
Figure 1. The animations in the first row are generated by a model with 1 billion parameters that ingests the entire audio recording. The animations in the second row are generated by our model $S_{5}+$, which has 0.8 million parameters (0.08%) and uses a context window of 512ms with a latency of 64ms (future length relative to the current frame). The audio was recorded with a headset microphone.
[…]
Teaser
@@ 18             Extract intermediate features: $\hat{f}=F^{S_{0}}(x)$ and $\tilde{f}=F^{S}(x)$;
19             Compute loss: $\mathcal{L}=\alpha_{y}\mathcal{L}(\hat{y},\tilde{y})+\alpha_{f}\mathcal{L}_{feat}(\hat{f},\tilde{f})$;
@@ 3.1.2. Student model
To predict animation $\mathbf{r}_{t}$ at frame $t$, the input to our student network is a window of raw waveform speech $\mathbf{w}^{d}_{t}$, sampled at $16$ kHz. We limit $\mathbf{w}^{d}_{t}$ to a $512$ ms segment of dimension $1\times 8192$. The audio window $\mathbf{w}^{d}_{t}$ consists of $512-d$ ms of past and $d$ ms future samples relative to the current frame $t$ (see Figure 2). This results in an initial latency of $d$ ms. We describe how to reduce this latency further in Section 3.2.
@@
Our student model compromises ten $1$D convolution layers, one linear layer, and three fully-connected layers, as illustrated in Table 1. This architecture is inspired by the first layers of HuBERT (Hsu et al., 2021) preceding the transformer layers.
@@
Conv 1D (×$\times 4$) 3 2 $C\times 101$ gelu
Conv 1D (×$\times 2$) 2 2 $C\times 25$ gelu
LN + Linear - - $C\times 25$
@@
LN + Conv 1D (×$\times 3$) 3 2 $C\times 2$ gelu
@@ 3.2.1. Down-scaling feature supervision
where $S$ and $S_{0}$ represent the student with fewer channels and the intermediate teacher, respectively. $F_{l}^{S}(\mathbf{w}_{t})$ is the feature map of layer $l$ corresponding to student $S$ and audio input $\mathbf{w}_{t}$, and $l=-n$ corresponds to the $n$-th-to-last layer in Table 1.
@@ 4.1.3. Implementation details
Our method is implemented using PyTorch (Paszke et al., 2019), and we use Adam (Kingma and Ba, 2015) to optimize the neural networks. We use a learning rate of $41e-$ 4 for training on LibriSpeech. Due to memory limitations, for each epoch we shuffle the training data and divide it into 48 subsets, training sequentially on each subset. Each student model is trained for 2–4 epochs, with each epoch taking approximately 24–48 hours on a single NVIDIA RTX 6000. For fine-tuning, we use a learning rate of 1e-6 for 10–20 epochs, with each epoch taking 1–2 minutes.
@@ 4.2.2. PBM accuracy
We evaluate lip-closure accuracy in animation frames corresponding to the bilabial consonants /p/, /b/, and /m/. Accurate lip closure during these sounds enhances the realism of the animation and is an important indicator of animation quality (Richard et al., 2021). Using HuBERT (Hsu et al., 2021) to label each frame’s phoneme, we efficiently select the relevant time stamps with minimal manual effort. The distance on the y-axis between the upper and lower lips is then calculated in mesh space, using a differentiable rig function  (Marquis Bolduc and Phan, 2022) (henceforth called rig2mesh) to convert rig parameters into mesh representations for each frame. We set the distance threshold for lip-closure to 0.15, as this is still visually perceived as a closed lip during animation in our case. To account for potential label misalignment introduced by HuBERT, we take the minimal distance value within ±$\pm 2$ frames of the current frame.
@@ 4.3.2. Quantitative metrics
$T$ vs $S_{0}$: $S_{0}$ achieves PBM accuracy comparable to the teacher model (99.25%) while using only 0.3% of the parameters, ¡0.8% of the FLOPs, and $\approx$0.5of the memory. This indicates that given a powerful enough student, we can achieve high-quality animations with significantly smaller models.
@@ 4.6.1. Receptive field of teacher
$\ \ T$ Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix A Details of Voice2Face teacher model
In the current Voice2Face model, we replace the original MFCC and SSC inputs with outputs from the penultimate layer of HuBERT. Previously, inputs had a shape of $[T,39]$, whereas the HuBERT features corresponding to an identical size time window have a shape of $[T^{\prime},1280]$ ($T^{\prime}<T$). To accommodate this change, we interpolate the HuBERT feature to $[T,1280]$ and modify the first convolution layer in the frequency processing network to handle the HuBERT feature dimension of 1280, ensuring the rest of the models in the pipeline remain unchanged.

### v2

@@ Abstract.
Project page: https://electronicarts.github.io/tiny-voice2face/
@@
submissionid: 1216journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4article: 104publicationmonth: 8doi: 10.1145/3730929ccs: Computing methodologies Neural networksccs: Computing methodologies Procedural animation
@@
Figure 1. The animations in the first row are generated by a model with 1 billion parameters that ingests the entire audio recording. The animations in the second row are generated by our model $S_{5}+$, which has 0.8 million parameters (0.08%) and uses a context window of 512ms with a latency of 64ms (future length relative to the current frame). The audio was recorded with a headset microphone.:
@@ 18      Extract intermediate features: $\hat{f}=F^{S_{0}}(x)$ and $\tilde{f}=F^{S}(x)$;
[…]
@@ 3.1.2. Student model
To predict animation $\mathbf{r}_{t}$ at frame $t$, the input to our student network is a window of raw waveform speech $\mathbf{w}^{d}_{t}$, sampled at $16$ kHz. We limit $\mathbf{w}^{d}_{t}$ to a $512$ ms segment of dimension $1\times 8192$. The audio window $\mathbf{w}^{d}_{t}$ consists of 512d512-d ms of past and $d$ ms future samples relative to the current frame $t$ (see Figure 2). This results in an initial latency of $d$ ms. We describe how to reduce this latency further in Section 3.2.
@@
Our student model compromises ten 11D convolution layers, one linear layer, and three fully-connected layers, as illustrated in Table 1. This architecture is inspired by the first layers of HuBERT (Hsu et al., 2021) preceding the transformer layers.
@@
Conv 1D ($\times 4$) 3 2 $C\times 101$ gelu
Conv 1D ($\times 2$) 2 2 $C\times 25$ gelu
LN + Linear - - $C\times 25$ -
@@
LN + Conv 1D ($\times 3$) 3 2 $C\times 2$ gelu
@@ 3.2.1. Down-scaling feature supervision
where $S$ and $S_{0}$ represent the student with fewer channels and the intermediate teacher, respectively. $F_{l}^{S}(\mathbf{w}_{t})$ is the feature map of layer $l$ corresponding to student $S$ and audio input $\mathbf{w}_{t}$, and l=nl=-n corresponds to the $n$-th-to-last layer in Table 1.
@@ 4.1.3. Implementation details
Our method is implemented using PyTorch (Paszke et al., 2019), and we use Adam (Kingma and Ba, 2015) to optimize the neural networks. We use a learning rate of 1e41e-4 for training on LibriSpeech. Due to memory limitations, for each epoch we shuffle the training data and divide it into 48 subsets, training sequentially on each subset. Each student model is trained for 2–4 epochs, with each epoch taking approximately 24–48 hours on a single NVIDIA RTX 6000. For fine-tuning, we use a learning rate of 1e-6 for 10–20 epochs, with each epoch taking 1–2 minutes.
@@ 4.2.2. PBM accuracy
We evaluate lip-closure accuracy in animation frames corresponding to the bilabial consonants /p/, /b/, and /m/. Accurate lip closure during these sounds enhances the realism of the animation and is an important indicator of animation quality (Richard et al., 2021). Using HuBERT (Hsu et al., 2021) to label each frame’s phoneme, we efficiently select the relevant time stamps with minimal manual effort. The distance on the y-axis between the upper and lower lips is then calculated in mesh space, using a differentiable rig function  (Marquis Bolduc and Phan, 2022) (henceforth called rig2mesh) to convert rig parameters into mesh representations for each frame. We set the distance threshold for lip-closure to 0.15, as this is still visually perceived as a closed lip during animation in our case. To account for potential label misalignment introduced by HuBERT, we take the minimal distance value within $\pm 2$ frames of the current frame.
@@ 4.3.2. Quantitative metrics
$T$ vs $S_{0}$: $S_{0}$ achieves PBM accuracy comparable to the teacher model (99.25%) while using only 0.3% of the parameters, ¡0.8% of the FLOPs, and $\approx$¡0.5% of the memory. This indicates that given a powerful enough student, we can achieve high-quality animations with significantly smaller models.
@@ 4.6.1. Receptive field of teacher
T\ \ T Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix A Details of Voice2Face teacher model
In the current Voice2Face model, we replace the original MFCC and SSC inputs with outputs from the penultimate layer of HuBERT. Previously, inputs had a shape of $[T,39]$, whereas the HuBERT features corresponding to an identical size time window have a shape of $T^{\prime},1280]$ ($T^{\prime}<T$). To accommodate this change, we interpolate the HuBERT feature to $[T,1280]$ and modify the first convolution layer in the frequency processing network to handle the HuBERT feature dimension of 1280, ensuring the rest of the models in the pipeline remain unchanged.